Refer to Interactive Solution 27.9 for help in solving this problem. In a Young's double-slit experiment the separation distance y between the second-order bright fringe and the central bright fringe on a flat screen is 0.0197 m, when the light has a wavelength of 425 nm. Assume that the angles are small enough so that sin is approximately equal to tan . Find the separation y when the light has a wavelength of 578 nm.
Answer #1

In Youngs double slit experiment ; the bright fringe is observed when y=mID (for small angle) & ya : Y2 = Si (2) :: 92 = 0·0